# Statistical Learning of Distributionally Robust Stochastic Control in Continuous State Spaces

### v1

@@ C.4 Proof of Theorem 3
$\bar{u}^{*}(x)$ $=\sup_{\phi\in\mathcal{Q}}\int_{\mathbb{A}}r(x,a)\phi(da)+\alpha\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}\int_{\mathbb{A}}\bar{u}^{*}(f(x,a,w))\phi(da)\psi(dw)$
$=\sup_{\phi\in\mathcal{Q}}\int_{\mathbb{A}}r(x,a)\phi(da)+\sup_{\lambda\geq 0}-\lambda\delta+\int_{\mathbb{W}}\inf_{y\in\mathbb{W}}\left[\int_{\mathbb{A}}\bar{u}^{*}(f(x,a,y))\phi(da)+\lambda c(w,y)\right]\mu_{0}(dw)$
By the same argument, the supremum is achieved within $\Lambda:=[0,\delta^{-1}\left\|\bar{u}^{*}\right\|]$. So, we have that
$\left\|\overline{\mathbf{T}}(\bar{u}^{*})(x)-\mathcal{T}(\bar{u}^{*})(x)\right\|$ $\leq\alpha\sup_{x\in\mathbb{X},\phi\in\mathcal{Q}}\left|\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{W}}\int_{\mathbb{A}}\bar{u}^{*}(f(x,a,w))\phi(da)\psi(dw)-\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}\int_{\mathbb{A}}\bar{u}^{*}(f(x,a,w))\phi(da)\psi(dw)\right|$
$\leq\alpha\sup_{x\in\mathbb{X},\phi\in\mathcal{Q},\lambda\in\Lambda}\left|\int_{\mathbb{W}}\inf_{y\in\mathbb{W}}\left[\int_{\mathbb{A}}\bar{u}^{*}(f(x,a,y))\phi(da)+\lambda c(w,y)\right](\mu_{0}-\hat{\mu})(dw)\right|$
@@
$\mathcal{G}:=\left\{{w\rightarrow\inf_{y\in\mathbb{W}}\left[\int_{\mathbb{A}}\bar{u}^{*}(f(x,a,y))\phi(da)+\lambda c(w,y)\right]:(x,\psi,\lambda)\in\Theta}\right\}.$
@@
$R_{n}(\mathbf{w},g):=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\epsilon_{i}g(w_{i})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\epsilon_{i}\inf_{y\in\mathbb{W}}\left[\int_{\mathbb{A}}\bar{u}^{*}(f(x,a,y))\phi(da)+\lambda c(w_{i},y)\right],$
@@
$\leq\exp\left(\frac{\xi^{2}}{2n}\sum_{i=1}^{n}\sup_{y\in\mathbb{W}}\left|\int_{\mathbb{A}}\bar{u}^{*}(f(x,a,w))\phi(da)-\int_{\mathbb{A}}\bar{u}^{*}(f(x^{\prime},a,w))\phi^{\prime}(da)+(\lambda-\lambda^{\prime})c(w_{i},y)\right|^{2}\right)$
$\leq\exp\left(\frac{\xi^{2}}{2}\left[\sup_{y\in\mathbb{W}}\left|\int_{\mathbb{A}}\bar{u}^{*}(f(x,a,y))\phi(da)-\int_{\mathbb{A}}\bar{u}^{*}(f(x^{\prime},a,y))\phi^{\prime}(da)\right|+c_{\vee}\left|\lambda-\lambda^{\prime}\right|\right]^{2}\right)$
@@
$\left|\int_{\mathbb{A}}\bar{u}^{*}(f(x,a,y))\phi(da)-\int_{\mathbb{A}}\bar{u}^{*}(f(x^{\prime},a,y))\phi^{\prime}(da)\right|$ (C.5)
$\leq\left|\int_{\mathbb{A}}\bar{u}^{*}(f(x,a,w))-u^{*}(f(x^{\prime},a,w))\phi(da)\right|+\left|\int_{\mathbb{A}}\bar{u}^{*}(f(x^{\prime},a,w))[\phi-\phi^{\prime}](da)\right|$
$\leq\ell|x-x^{\prime}|+\min\left\{{\beta\left\|\phi-\phi^{\prime}\right\|_{\mathrm{TV}},\ell W_{1}(\phi,\phi^{\prime})}\right\}$
@@
$\rho(\theta,\theta^{\prime}):=\ell|x-x^{\prime}|+\beta\left\|\phi-\phi^{\prime}\right\|_{\mathrm{TV}}+c_{\vee}|\lambda-\lambda^{\prime}|$
@@
Therefore, using Dudley’s entropy integral [Wainwright, 2019, Chapter 5], the empirical Rademacher complexity can be bounded
[…]
@@
$\overline{D}_{\vee}:=\ell\mathrm{diam}(\mathbb{X})+2\beta+c_{\vee}\delta^{-1}\beta+1\geq\sup_{g,g^{\prime}\in\mathcal{G}}\rho(g,g^{\prime}).$
@@
[…]
@@
$\mathcal{N}(\epsilon;\Theta,\rho)$ $\leq\mathcal{N}(\epsilon;\mathbb{X},\ell|\cdot|)\cdot\mathcal{N}(\epsilon;\mathcal{Q},\beta\left\|\cdot\right\|_{\mathrm{TV}})\cdot\mathcal{N}(\epsilon;\Lambda,c_{\vee}|\cdot|)$ (C.6)
$\leq\mathcal{N}(\epsilon;\mathbb{X},\ell|\cdot|)\cdot\mathcal{N}(\epsilon;B_{1}^{|\mathbb{A}|},\beta\left\|\cdot\right\|_{1})\cdot\mathcal{N}(\epsilon;\Lambda,c_{\vee}|\cdot|)$
$\leq\left(1+\frac{\ell\mathrm{diam}(\mathbb{X})}{\epsilon}\right)^{d_{\mathbb{X}}}\left(1+\frac{2\beta}{\epsilon}\right)^{|\mathbb{A}|}\left(\frac{c_{\vee}\mathrm{diam}(\Lambda)}{\epsilon}+1\right).$
where $B_{1}^{|\mathbb{A}|}$ is the $|\mathbb{A}|$ dimensional $\ell_{1}$-ball of radius 1, and its covering number bound follows from Wainwright [2019, Example 5.8]. Therefore, the entropy integral
[…]
@@
$\stackrel{{\scriptstyle(i)}}{{\leq}}2\alpha\beta\mathcal{R}_{n}(\mathcal{G})+\alpha\beta\sqrt{\frac{2g_{\vee}}{n}\log{\frac{1}{\eta}}}$
$\leq\left(32\sqrt{\pi}\alpha\beta\overline{D}_{\vee}\sqrt{d_{\mathbb{X}}+|\mathbb{A}|+1}+\sqrt{2}\alpha\beta^{3/2}\sqrt{2\log{\frac{1}{\eta}}}\right)n^{-1/2}$
where $(i)$ holds w.p. at least $1-\eta$.
@@ C.5 Proof of Theorem 4
$\left\|\overline{\mathbf{T}}(\bar{u}^{*})(x)-\mathcal{T}(\bar{u}^{*})(x)\right\|$ $\leq\alpha\sup_{x\in\mathbb{X},\phi\in\mathcal{Q}}\left|\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{W}}\int_{\mathbb{A}}\bar{u}^{*}(f(x,a,w))\phi(da)\psi(dw)-\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}\int_{\mathbb{A}}\bar{u}^{*}(f(x,a,w))\phi(da)\psi(dw)\right|$
$=\alpha\sup_{\theta\in\Theta}\left|\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{W}}g_{\theta}(w)\psi(dw)-\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}g_{\theta}(w)\psi(dw)\right|$
@@
$g_{\theta}(w):=\int_{\mathbb{A}}\bar{u}^{*}(f(x,a,w))\phi(da).$
By Duchi and Namkoong [2021, Corollary 1], for $n\geq k\vee 3$, the r.h.s. satisfies w.p. at least $1-2\mathcal{N}(\epsilon/3;\mathcal{G},\left\|\cdot\right\|)e^{-t}$,
@@
$\epsilon=n^{-\frac{1}{k^{\prime}\vee 2}}c_{k}(\rho)^{2}\left(\frac{c_{k}(\rho)}{c_{k}(\rho)-1}\vee 2\right)\left(\frac{1}{k}+\sqrt{t+2\log n}\right).$
@@
$\left\|\mathbf{\overline{T}}(\bar{u}^{*})-\overline{\mathcal{T}}(\bar{u}^{*})\right\|\leq 30\alpha\beta n^{-\frac{1}{k^{\prime}\vee 2}}c_{k}(\rho)^{2}\left(\frac{c_{k}(\rho)}{c_{k}(\rho)-1}\vee 2\right)\left(\frac{1}{k}+\sqrt{\log(2\mathcal{N}(\epsilon/3;\mathcal{G},\left\|\cdot\right\|)+\log\frac{1}{\eta}+2\log n}\right).$
@@
$d((x,\phi),(x^{\prime},\phi^{\prime}))=\ell|x-x^{\prime}|+\beta\left\|\phi-\phi^{\prime}\right\|_{\mathrm{TV}}.$
This and the Lipschitz covering number bound [van der Vaart and Wellner, 1996, Chapter 2.7.4] implies that
$\log\mathcal{N}(\epsilon/3;\Theta,\rho)$ $\leq d_{\mathbb{X}}\log\left(1+\frac{3\ell\mathrm{diam}(\mathbb{X})}{\epsilon\wedge 1}\right)+|\mathbb{A}|\log\left(1+\frac{6\beta}{\epsilon\wedge 1}\right)$
$\leq d_{\mathbb{X}}\log\left(1+3\ell\mathrm{diam}(\mathbb{X})\right)+|\mathbb{A}|\log\left(1+6\beta\right)+\frac{1}{2}\left(d_{\mathbb{X}}+|A|\right)\log n$
@@
Therefore, defining $\overline{D}:=d_{\mathbb{X}}\log\left(1+3\ell\mathrm{diam}(\mathbb{X})\right)+|\mathbb{A}|\log\left(1+6\beta\right)$, we conclude that
$\left\|\hat{u}-u^{*}\right\|$ $\leq 30\beta^{2}n^{-\frac{1}{k^{\prime}\vee 2}}c_{k}(\rho)^{2}\left(\frac{c_{k}(\rho)}{c_{k}(\rho)-1}\vee 2\right)\left(\frac{1}{k}+\sqrt{\overline{D}+\log\frac{1}{\eta}+2(d_{\mathbb{X}}+|\mathbb{A}|)\log n}\right)$
@@ D.1 Proof of Lemma 1
which trivially lower bounds the uniform version. We will prove Theorem 5 and 6 by showing the same lower bound for this local risk.
@@ D.3 Proof of Theorem 6
Here, $\beta(1-Z)$ has a two-point distribution on $\left\{{0,\beta}\right\}$ under $\mu$. Therefore, Theorem 3 of Duchi and Namkoong [2021] applies. Define

### v2

@@ C.4 Proof of Theorem 4
$\bar{u}^{*}(x)$ $=\sup_{\phi\in\mathcal{Q}}\int_{\mathbb{A}}\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}[r(x,a,w)+\alpha\bar{u}^{*}(f(x,a,w))]\psi(dw)\phi(da)$
$=\sup_{\phi\in\mathcal{Q}}\int_{\mathbb{A}}\sup_{\lambda\geq 0}-\lambda\delta+\int_{\mathbb{W}}\inf_{y\in\mathbb{W}}\left[[r(x,a,y)+\alpha\bar{u}^{*}(f(x,a,y))]+\lambda c(w,y)\right]\mu_{0}(dw)\phi(da)$
By the same argument, the supremum is achieved within $\Lambda:=[0,\delta^{-1}(1+\alpha\left\|\bar{u}^{*}\right\|)]$. So, we have that
$\left\|\overline{\mathbf{T}}(\bar{u}^{*})(x)-\mathcal{T}(\bar{u}^{*})(x)\right\|$ $\leq\sup_{x\in\mathbb{X},\phi\in\mathcal{Q}}\left|\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{W}}\int_{\mathbb{A}}[r(x,a,w)+\alpha\bar{u}^{*}(f(x,a,w))]\phi(da)\psi(dw)-\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}\int_{\mathbb{A}}[r(x,a,w)+\alpha\bar{u}^{*}(f(x,a,w))]\phi(da)\psi(dw)\right|$
$\leq\sup_{x\in\mathbb{X},\phi\in\mathcal{Q},\lambda\in\Lambda}\left|\int_{\mathbb{W}}\inf_{y\in\mathbb{W}}\left[\int_{\mathbb{A}}[r(x,a,y)+\alpha\bar{u}^{*}(f(x,a,y))]\phi(da)+\lambda c(w,y)\right](\mu_{0}-\hat{\mu})(dw)\right|$
@@
$\mathcal{G}:=\left\{{w\rightarrow\inf_{y\in\mathbb{W}}\left[\int_{\mathbb{A}}[r(x,a,y)+\alpha\bar{u}^{*}(f(x,a,y))]\phi(da)+\lambda c(w,y)\right]:(x,\psi,\lambda)\in\Theta}\right\}.$
@@
$R_{n}(\mathbf{w},g):=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\epsilon_{i}g(w_{i})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\epsilon_{i}\inf_{y\in\mathbb{W}}\left[\int_{\mathbb{A}}[r(x,a,y)+\alpha\bar{u}^{*}(f(x,a,y))]\phi(da)+\lambda c(w_{i},y)\right],$
@@
$\leq\exp\left(\frac{\xi^{2}}{2n}\sum_{i=1}^{n}\sup_{y\in\mathbb{W}}\left|\int_{\mathbb{A}}[r(x,a,y)+\alpha\bar{u}^{*}(f(x,a,y))]\phi(da)-\int_{\mathbb{A}}[r(x^{\prime},a,y)+\alpha\bar{u}^{*}(f(x^{\prime},a,y))]\phi^{\prime}(da)+(\lambda-\lambda^{\prime})c(w_{i},y)\right|^{2}\right)$
$\leq\exp\left(\frac{\xi^{2}}{2}\left[\sup_{y\in\mathbb{W}}\left|\int_{\mathbb{A}}[r(x,a,y)+\alpha\bar{u}^{*}(f(x,a,y))]\phi(da)-\int_{\mathbb{A}}\bar{[}r(x^{\prime},a,y)+\alpha\bar{u}^{*}(f(x^{\prime},a,y))]\phi^{\prime}(da)\right|+c_{\vee}\left|\lambda-\lambda^{\prime}\right|\right]^{2}\right)$
@@
$\left|\int_{\mathbb{A}}[r(x,a,y)+\alpha\bar{u}^{*}(f(x,a,y))]\phi(da)-\int_{\mathbb{A}}[r(x^{\prime},a,y)+\alpha\bar{u}^{*}(f(x^{\prime},a,y))]\phi^{\prime}(da)\right|$ (C.5)
$\leq\left|\int_{\mathbb{A}}[r(x,a,y)+\alpha\bar{u}^{*}(f(x,a,y))]-[r(x^{\prime},a,y)+\alpha\bar{u}^{*}(f(x^{\prime},a,y))]\phi(da)\right|+\left|\int_{\mathbb{A}}[r(x^{\prime},a,y)+\alpha\bar{u}^{*}(f(x^{\prime},a,y))][\phi-\phi^{\prime}](da)\right|$
$\leq L|x-x^{\prime}|+\min\left\{{\beta\left\|\phi-\phi^{\prime}\right\|_{\mathrm{TV}},LW_{1}(\phi,\phi^{\prime})}\right\}$
@@
$\rho(\theta,\theta^{\prime}):=L|x-x^{\prime}|+\beta\left\|\phi-\phi^{\prime}\right\|_{\mathrm{TV}}+c_{\vee}|\lambda-\lambda^{\prime}|$
@@
Therefore, using Dudley’s entropy integral [45, Chapter 5], the empirical Rademacher complexity can be bounded
$\mathcal{R}_{n}(\mathbf{w},\mathcal{G})\leq\frac{32}{\sqrt{n}}\int_{0}^{D_{\vee}}\sqrt{\log\mathcal{N}(\epsilon;\Theta,\rho)}d\epsilon$
@@
$\overline{D}_{\vee}:=L\mathrm{diam}(\mathbb{X})+2\beta+c_{\vee}\delta^{-1}\beta\geq\sup_{g,g^{\prime}\in\mathcal{G}}\rho(g,g^{\prime}).$
@@
$\mathcal{R}_{n}(\mathcal{G})\leq\frac{32}{\sqrt{n}}\int_{0}^{D_{\vee}}\sqrt{\log\mathcal{N}(\epsilon;\Theta,\rho)}d\epsilon$
@@
$\mathcal{N}(\epsilon;\Theta,\rho)$ $\leq\mathcal{N}(\epsilon;\mathbb{X},L|\cdot|)\cdot\mathcal{N}(\epsilon;\mathcal{Q},\beta\left\|\cdot\right\|_{\mathrm{TV}})\cdot\mathcal{N}(\epsilon;\Lambda,c_{\vee}|\cdot|)$ (C.6)
$\leq\mathcal{N}(\epsilon;\mathbb{X},L|\cdot|)\cdot\mathcal{N}(\epsilon;B_{1}^{|\mathbb{A}|},\beta\left\|\cdot\right\|_{1})\cdot\mathcal{N}(\epsilon;\Lambda,c_{\vee}|\cdot|)$
$\leq\left(1+\frac{L\mathrm{diam}(\mathbb{X})}{\epsilon}\right)^{d_{\mathbb{X}}}\left(1+\frac{2\beta}{\epsilon}\right)^{|\mathbb{A}|}\left(\frac{c_{\vee}\mathrm{diam}(\Lambda)}{\epsilon}+1\right).$
where $B_{1}^{|\mathbb{A}|}$ is the $|\mathbb{A}|$ dimensional $\ell_{1}$-ball of radius 1, and its covering number bound follows from Wainwright, [45, Example 5.8]. Therefore, the entropy integral
$\int_{0}^{\overline{D}_{\vee}}\sqrt{\log\mathcal{N}(\epsilon;\Theta,\rho)}d\epsilon$ $\leq\int_{0}^{\overline{D}_{\vee}}\sqrt{(d_{\mathbb{X}}+|\mathbb{A}|+1)\log(\overline{D}_{\vee}/\epsilon)}d\epsilon$
@@
$\stackrel{{\scriptstyle(i)}}{{\leq}}2\beta\mathcal{R}_{n}(\mathcal{G})+\beta\sqrt{\frac{2g_{\vee}}{n}\log{\frac{1}{\eta}}}$
$\leq\left(32\sqrt{\pi}\beta\overline{D}_{\vee}\sqrt{d_{\mathbb{X}}+|\mathbb{A}|+1}+\sqrt{2}\beta^{3/2}\sqrt{2\log{\frac{1}{\eta}}}\right)n^{-1/2}$
where $(i)$ holds w.p. at least $1-\eta$. This bound implies the theorem as we note that $d_{\mathbb{X}}\geq 1$.
@@ C.5 Proof of Theorem 5
$\left\|\overline{\mathbf{T}}(\bar{u}^{*})(x)-\mathcal{T}(\bar{u}^{*})(x)\right\|$ $\leq\sup_{x\in\mathbb{X},\phi\in\mathcal{Q}}\left|\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{W}}\int_{\mathbb{A}}[r(x,a,w)+\alpha\bar{u}^{*}(f(x,a,w))]\phi(da)\psi(dw)-\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}\int_{\mathbb{A}}[r(x,a,w)+\alpha\bar{u}^{*}(f(x,a,w))\phi(da)\psi(dw)\right|$
$=\sup_{\theta\in\Theta}\left|\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{W}}g_{\theta}(w)\psi(dw)-\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}g_{\theta}(w)\psi(dw)\right|$
@@
$g_{\theta}(w):=\int_{\mathbb{A}}[r(x,a,w)+\alpha\bar{u}^{*}(f(x,a,w))]\phi(da).$
By Duchi and Namkoong, [9, Corollary 1], for $n\geq k\vee 3$, the r.h.s. satisfies w.p. at least $1-2\mathcal{N}(\epsilon/3;\mathcal{G},\left\|\cdot\right\|)e^{-t}$,
@@
$\epsilon=n^{-\frac{1}{k^{\prime}\vee 2}}c_{k}(\delta)^{2}\left(\frac{c_{k}(\delta)}{c_{k}(\delta)-1}\vee 2\right)\left(\frac{1}{k}+\sqrt{t+2\log n}\right).$
@@
$\left\|\mathbf{\overline{T}}(\bar{u}^{*})-\overline{\mathcal{T}}(\bar{u}^{*})\right\|\leq 30\beta n^{-\frac{1}{k^{\prime}\vee 2}}c_{k}(\delta)^{2}\left(\frac{c_{k}(\delta)}{c_{k}(\delta)-1}\vee 2\right)\left(\frac{1}{k}+\sqrt{\log(2\mathcal{N}(\epsilon/3;\mathcal{G},\left\|\cdot\right\|)+\log\frac{1}{\eta}+2\log n}\right).$
@@
$\rho((x,\phi),(x^{\prime},\phi^{\prime}))=L|x-x^{\prime}|+\beta\left\|\phi-\phi^{\prime}\right\|_{\mathrm{TV}}.$
This and the Lipschitz covering number bound [43, Chapter 2.7.4] implies that
$\log\mathcal{N}(\epsilon/3;\Theta,\rho)$ $\leq d_{\mathbb{X}}\log\left(1+\frac{3L\mathrm{diam}(\mathbb{X})}{\epsilon\wedge 1}\right)+|\mathbb{A}|\log\left(1+\frac{6\beta}{\epsilon\wedge 1}\right)$
$\leq d_{\mathbb{X}}\log\left(1+3L\mathrm{diam}(\mathbb{X})\right)+|\mathbb{A}|\log\left(1+6\beta\right)+\frac{1}{2}\left(d_{\mathbb{X}}+|A|\right)\log n$
@@
Therefore, defining $\overline{D}:=d_{\mathbb{X}}\log\left(1+3L\mathrm{diam}(\mathbb{X})\right)+|\mathbb{A}|\log\left(1+6\beta\right)$, we conclude that
[…]
@@ D.1 Proof of Lemma 1
which trivially lower bounds the uniform version. We will prove Theorem 6 and 7 by showing the same lower bound for this local risk.
@@ D.3 Proof of Theorem 7
Here, $\beta(1-Z)$ has a two-point distribution on $\left\{{0,\beta}\right\}$ under $\mu$. Therefore, Theorem 3 of Duchi and Namkoong, [9] applies. Define
@@
Appendix E Algorithm Design for the CAU Case
To parameterize randomized controller policies, we employ a generative model (Algorithm 2) by considering $\pi_{\eta}:\mathbb{R}^{d}\times\mathbb{X}\rightarrow\mathbb{A}$ where an action is generated by $A\sim\pi_{\eta}(N,x)$ using an independence source of randomness $N\sim N(0,I)$, a standard Gaussian vector independent of the state.
Under this definition, we overwrite the notation and consider
$\widehat{\mathbf{T}}_{\eta,\theta}(\lambda,x):=\lambda-c_{k}(\delta)\left[\int_{\mathbb{W}}(E_{N}[r(x,\pi_{\eta}(N,x))+\alpha u_{\theta}(f(x,\pi_{\eta}(N,x),w))]-\lambda)_{+}^{k^{\prime}}\hat{\mu}(dw)\right]^{1/k^{\prime}}$
By strong duality, the Bellman operator under policy $\pi_{\eta}$ applied to $u_{\theta}$ is
$\sup_{\lambda\in\mathbb{R}}\left(\lambda-c_{k}(\delta)\left[\int_{\mathbb{W}}\left(E_{N}[r(x,\pi_{\eta}(N,x),w)+\alpha u_{\theta}(f(x,\pi_{\eta}(N,x),w))]-\lambda\right)^{k^{\prime}}\hat{\mu}(dw)\right]^{1/k^{\prime}}\right)=\widehat{\mathbf{T}}_{\eta,\theta}(\lambda^{\ast},x)$
where $\lambda^{\ast}=\lambda^{\ast}(\eta,\theta,x)$ is the optimal dual multiplier. We note that $\lambda^{\ast}$ doesn’t depend on the realizations of $N$, and can be computed via bisection search.
Bellman Error Minimization: Analogous to the CAA case, we minimize the $L^{2}$ Bellman error:
$\min_{\theta}\int_{\mathbb{X}}[u_{\theta}(x)-\widehat{\mathbf{T}}_{\eta,\theta}(\lambda^{\ast},x)]^{2}\nu(dx)$
The gradient is evaluated using the envelope theorem
$\nabla_{\theta}\mathbf{T}_{\eta,\theta}(x)=\nabla_{\theta}\widehat{\mathbf{T}}_{\eta,\theta}(\lambda^{\ast}(\eta,\theta,x),x)$
and the expectation over $N$ can be approximated using the sample average over $m$ i.i.d. samples $N_{1},\dots,N_{m}$.
Therefore, we update the $\theta$ using first order methods; e.g. mini-batch stochastic gradient descent (SGD):
$\theta_{t+1}=\theta_{t}-\ell_{t}G_{n,m,t}$
where $\ell_{t}$ is the learning rate and $G_{n,m,t}$ is a stochastic gradient.
The gradient estimate $G_{n,m,t}$ can be obtained as follows: we first sample $\mathbf{N}=\left\{{N_{1},\dots,N_{m}}\right\}$ and $X_{i}\sim\nu$ i.i.d. and compute $\lambda_{m,i}^{\ast}=\lambda^{\ast}(\eta,\theta,\mathbf{N},X_{i})$ that maximize
$\lambda-c_{k}(\delta)\left[\int_{\mathbb{W}}\left(\frac{1}{m}\sum_{j=1}^{m}u_{\theta}(f(X_{i},\pi_{\eta}(N_{j},X_{i}),w))-\lambda\right)^{k^{\prime}}\hat{\mu}(dw)\right]^{1/k^{\prime}}$
using bisection search. Then, we set
$G_{n,m,t}=\frac{1}{n}\sum_{i=1}^{n}2(u_{\theta}(X_{i})-\widehat{\mathbf{T}}_{\eta,\theta}(\lambda_{m,i}^{\ast},X_{i}))(\nabla_{\theta}u_{\theta}(X_{i})-\nabla_{\theta}\overline{\mathbf{T}}_{\eta,\theta}(\lambda^{\ast}_{m,i},X_{i})).$
Policy Improvement: The policy improvement step parallels the CAA setting:
$\eta_{t+1}=\eta_{t}+\ell^{\prime}_{t}\frac{1}{n}\sum_{i=1}^{n}\nabla_{\eta}\widehat{\mathbf{T}}_{\eta,\theta}(\lambda^{\ast}_{m,i},X_{i}),$
for some possibly different learning rate $\ell^{\prime}_{t}$.
Algorithm 2 Robust Actor–Critic Algorithm for CAU
1:Step–size schedules $\{\ell_{t}\}_{t\geq 0},\{\ell^{\prime}_{t}\}_{t\geq 0}$; state–batch size $n$; noise–sample size $m$; initial parameters $(\theta_{0},\eta_{0})$; sampling measure $\nu$ on $\mathbb{X}$
2:Repeat until convergence:
3:
4:1. Bellman–error (critic) update
5:Sample states $\{X_{i}\}_{i=1}^{n}\overset{\text{i.i.d.}}{\sim}\nu$
6:For each $i$, draw i.i.d. noise vectors $\{N_{ij}\}_{j=1}^{m}\sim\mathcal{N}(0,I_{d})$
7:for $i=1,\dots,n$ in parallel do
8:  $\lambda_{i}^{\!*}\;\leftarrow\;\arg\max_{\lambda\in\mathbb{R}}\;\widehat{\mathbf{T}}_{\eta_{t},\theta_{t}}\!\bigl(\lambda,X_{i};\{N_{ij}\}_{j=1}^{m}\bigr)$ $\triangleright$ bisection search
9:end for
[…]
12:2. Policy–improvement (actor) update
13:Resample (or reuse) $\{X_{i}\},\{N_{ij}\}$
14:for $i=1,\dots,n$ in parallel do
15:  Recompute $\lambda_{i}^{\!*}\leftarrow\arg\max_{\lambda}\widehat{\mathbf{T}}_{\eta_{t},\theta_{t+1}}(\lambda,X_{i};\{N_{ij}\})$
16:end for
[…]
20:  return final parameters $(\theta_{t},\eta_{t})$
Appendix F Additional Experiment Results and Implementation Details
F.1 Inventory Control
F.1.1 Implementation Details
Actor Network
We use a 3-layer neural network with ReLU as the activation function, 8 neurons in each hidden layer. We treat $X_{t}$ as the CAA input, and $A_{t}$ as the output. With CAU, we have another standard normal random variable $N_{1}$ as the input.
Critic Network
We use a 3-layer neural network with ReLU as the activation function, 32 neurons in each hidden layer. We treat $X_{t}$ as the input, and the value of the policy $u(X_{t})$ as the output.
Training
To achieve convergence of the actor-critic network, we choose a small batch size (1), a small learning rate ($3\times 10^{-4}$, Adam optimizer) and a large number of epochs ($5\times 10^{4}$). For each of CAA and CAU cases, we train 5 sets of robust value and policy under $\delta\in\{1\times 10^{-3},1\times 10^{-2},1\times 10^{-1},5\times 10^{-1},1\}$.
F.1.2 Discussions on Numerical Techniques
Rescale the Data
We rescale the demand of the good by multiplying $10^{-3}$ to the original data. This results in faster convergence of the robust value and policy because we could have a smaller sample space for $X_{t}$ than before. As a consequence, our reported test result should be multiplied by $10^{3}$.
Compact Sample Space
Similar as the previous case, we need a compact support for $\nu(x)$ to ensure convergence. In this case, we observe that all the demand data lies in $[0,10]$ after rescaling. Thus, we choose $\nu(x)$ to be uniform on $[0,10]$.
Imposing i.i.d. Data Assumption
In reality, the demand data may not be i.i.d. as we assume here. For example, the demand would surge during holidays and may have a seasonal pattern. To make the demand data as i.i.d. as possible, we only keep the data with features Open = 1, Promo=0, StateHoliday=0, SchoolHoliday=0, namely, the regular demand not during promotion or holiday.
(a) CAA policies for Store 1
(b) CAU policies for Store 1
Figure 4: Robust policies for one store in the inventory control problem.
F.2 Portfolio Optimization
F.2.1 Implementation Details
Actor Network
We use a 3-layer neural network with ReLU as the activation function, 8 neurons in each hidden layer. For simplicity, we treat $\mathbf{1}^{\top}X_{t}$ as the CAA input, and $A_{t}^{\top}X_{t}$ as the output, rather than acquire a percentage $A_{t}$ first. With CAU, we have another two standard normal random variables $N_{1}$ and $N_{2}$ as the input.
Critic Network
We use a 3-layer neural network with ReLU as the activation function, 32 neurons in each hidden layer. We treat $\mathbf{1}^{\top}X_{t}$ as the input, and $u(\mathbf{1}^{\top}X_{t})$ as the output.
Training
To achieve convergence of the actor-critic network, we choose a small batch size (1), a small learning rate ($3\times 10^{-5}$, Adam optimizer) and a large number of epochs ($10^{5}$). For each of CAA and CAU cases, we train 7 sets of robust value and policy under $\delta\in\{5\times 10^{-4},1\times 10^{-3},5\times 10^{-3},1\times 10^{-2},5\times 10^{-2},1\times 10^{-1},5\times 10^{-1}\}$.
F.2.2 Discussions on Numerical Techniques
Compact Sample Space
Note that in our previous discussion, we should sample from $\nu(x)$ which is supported on $\mathbb{R}^{2}$. However, this would result in very slow convergence or no convergence at all. As a consequence, we choose a compact sample space $\{(x_{1},x_{2}):x_{1}+x_{2}\geq 0,x_{1}\leq 100,x_{2}\leq 100\}$. While this adjustment boosts convergence, it also induces bias in the robust value estimation. Empirically, we find the bias small and negligible.
Enforcing Uniform Lipschitz Condition
In Section 4 we assume the uniform Lipschitzness of the value function. However, in our setting, this is not the case in a small neighborhood around $\mathbf{1}^{\top}X_{t}=0$. Based on the form of the reward $r(X_{t},A_{t},W_{t})=\frac{C_{t}^{1/2}}{1/2}$, we expect the robust value $u(\mathbf{1}^{\top}X_{t})$ to behave like a square-root function. As a result, in order to enforce the uniform Lipschitz condition, we make the following change to the critic network: in the output layer, we multiply the output by $\sqrt{\mathbf{1}^{\top}X_{t}}$. This essentially results in a re-parametrized robust value function
$u_{\theta}(x)=v_{\theta}(x)\sqrt{\mathbf{1}^{\top}x},$
where $v_{\theta}(x)$ is the output of the critic network.
This makes the output of the critic network look like a constant function roughly. We find this technique particularly helpful with fitting the robust value and avoiding gradient explosion around $\mathbf{1}^{\top}x=0$, where the Lipschitz condition is violated.
Action Clipping
In order to enforce the two constraints mentioned above: $A_{t}^{\top}X_{t}\geq 0$ and $A_{t}^{\top}X_{t}\leq\mathbf{1}^{\top}X_{t}$, respectively, we clip the output of the actor network. If $A_{t}^{\top}X_{t}<0$, then we make it to 0 by adding each output by $\frac{-A_{t}^{\top}X_{t}}{m}$; if $A_{t}^{\top}X_{t}>\mathbf{1}^{\top}X_{t}$, then we make it to $\mathbf{1}^{\top}X_{t}$ by decreasing each action by $\frac{A_{t}^{\top}X_{t}-\mathbf{1}^{\top}X_{t}}{m}$.